OpenStudy (mathmath333):

fill in the boxes \(\huge \tt \begin{align} \color{black}{ \fbox{-}+ \fbox{-}+ \fbox{-}+ \fbox{-}+ \fbox{-}= \fbox{30}}\end{align}\) u can use \(\large\tt \begin{align} \color{black}{\{1,3,5,7,9,11,13,15\}}\end{align}\) u can repeat the numbers

ganeshie8 (ganeshie8):

cool problem and latex xD

OpenStudy (mathmath333):

lol

OpenStudy (anonymous):

test it with the largest number first

OpenStudy (anonymous):

interesting

OpenStudy (anonymous):

(15-9)+..that kind of thing?

OpenStudy (anonymous):

or actually can we use factorials?

OpenStudy (anonymous):

I can see it with factorials easy haha

OpenStudy (mathmath333):

no only one or two digit number

OpenStudy (anonymous):

mhmmmmm

OpenStudy (mathmath333):

yes but how ?

OpenStudy (anonymous):

i give u a rule

OpenStudy (anonymous):

I don't understand I don't think adding 5 odd numbers will give even

OpenStudy (anonymous):

i tried lol

OpenStudy (anonymous):

u have 5 places your answer is30 so in each of them should be 6 ok?

OpenStudy (anonymous):

Impossible given the numbers, unless there's a trick to this

OpenStudy (anonymous):

for example u can choose 9&3 instead of 6&6 then u can find it

OpenStudy (anonymous):

there has to be some extra rules

ganeshie8 (ganeshie8):

if there no solution, prove it @iambatman

OpenStudy (anonymous):

it has no answer be sure

OpenStudy (anonymous):

because u have 5 places with unit numbers

OpenStudy (anonymous):

and u have an even number in your answer

ganeshie8 (ganeshie8):

I hope rotating the numbers is not allowed :P \[\huge \tt \begin{align} \color{black}{ \fbox{6}+ \fbox{6}+ \fbox{6}+ \fbox{6}+ \fbox{6}= \fbox{30}}\end{align}\]

OpenStudy (mathmath333):

cool

OpenStudy (anonymous):

Actually no what I said was right lol

OpenStudy (anonymous):

There is no answer

OpenStudy (anonymous):

exactly iambatman

ganeshie8 (ganeshie8):

more mathematically \[\large \tt \begin{align} \color{black}{ \fbox{odd}+ \fbox{odd}+ \fbox{odd}+ \fbox{odd}+ \fbox{odd} \stackrel{?}{=} \fbox{30}}\end{align}\] \[\large \tt \begin{align} \color{black}{ ( \fbox{odd}+ \fbox{odd})+ (\fbox{odd}+ \fbox{odd})+ \fbox{odd} \stackrel{?}{=} \fbox{30}}\end{align}\] \[\large \tt \begin{align} \color{black}{ ( even )+ (even )+ \fbox{odd} \stackrel{?}{=} \fbox{30}}\end{align}\] \[\large \tt \begin{align} \color{black}{ (even )+ \fbox{odd} \stackrel{?}{=} \fbox{30}}\end{align}\]

OpenStudy (anonymous):

Nice haha.

OpenStudy (anonymous):

You could quickly just find the average and realize right away without having to do all this stuff haha xD

OpenStudy (mathmath333):

also this way \(\Large \tt \begin{align} \color{black}{5(2n-1) (mod 2) \\ =10n-5) (mod 2) \equiv 1}\end{align}\) its odd where \(\large n\in \mathbb{z} \) so no solution

ganeshie8 (ganeshie8):

nice :) but there is a small catch in your argument you're assuming all numbers are same

OpenStudy (anonymous):

Checkmate!

OpenStudy (mathmath333):

every odd positive integer can be represented as \((2n-1)\) ?

OpenStudy (anonymous):

Yes

OpenStudy (mathmath333):

so i think it includes all odd numbers

ganeshie8 (ganeshie8):

5(2n-1) does not represent sum of "5 arbitrary odd numbers" it represents adding "one arbitrary" odd number 5 times

OpenStudy (mathmath333):

ok this should be it \(\large\tt \begin{align} \color{black}{\left( (2n_1-1) (mod~ 2)+(2n_2-1) (mod ~2)\\+(2n_3-1) (mod~ 2)+(2n_4-1) (mod ~2) \\+(2n_5-1) (mod ~2) \\ \right)(mod~ 2) \equiv 1}\end{align}\) \(\{n_1,n_2,n_3,n_4,n_4,n_5\}\subset \mathbb{Z}\)